OpenStudy (anonymous):

find the particular solution that satisfies the initial conditions f''(x)= sinx + e^2x f(0)= 1/4, f'(0) =1/2

OpenStudy (accessdenied):

How much have you done with this problem so far?

OpenStudy (accessdenied):

Or haven't found the right place to begin? :)

rvc (rvc):

@pnp97 what exactly u want?

OpenStudy (anonymous):

i was sick for almost the whole chapter so i have no idea how to begin @rvc @AccessDenied

rvc (rvc):

derivative of sinx?

OpenStudy (accessdenied):

Hm.. alright. We are given the second derivative of a function. Our goal is to find the original function given that alone and some initial conditions (basically, just points that satisfy the equation). To get back from the derivative, we need integration.

OpenStudy (anonymous):

so take the integral of sinx+ e^2x??

OpenStudy (accessdenied):

Yep. \( \dfrac{d^2}{dx^2} f(x) = \sin x + e^{2x} \) \( \dfrac{d}{dx} \left( \dfrac{df}{dx} \right) = \sin x + e^{2x} \) We can integrate both sides with respect to x to cancel the first derivative.

OpenStudy (anonymous):

integral of sinx is -cosx and would e^2x stay the same??

OpenStudy (kainui):

As an example I just made up on the spot: conditions: f(1)=5 f'(0)=-2 differential equation: f''(x)=2x So integrate both sides you get: f'(x)=x^2+C since f'(0)=-2 then -2=C f'(x)=x^2-2 integrate again! \[f(x)=\frac{ 1 }{ 3 }x^3-2x+C\] Since f(1)=5 \[5=\frac{ 1 }{ 3 }(1)^3-2(1)+C\] solve for the constant again and then plug it in and you get your final, particular equation of f((x). if you didn't have the conditions you wouldn't have solved for the constants so you'd have C1 and C2 which would be the General solution rather than the Particular one.

OpenStudy (accessdenied):

Almost. If you took the derivative, d/dx e^(2x) = 2 e^(2x). You'll need to divide by 2 when integrating, because then by chain rule it works out when taking the derivative again as a check .

OpenStudy (anonymous):

so you have to integrate twice? @kainui

OpenStudy (anonymous):

ah so e^2x /2

OpenStudy (accessdenied):

Yes. Two integrations are required in total, this process is the first of the two, which we will add a constant C. And correct! :)

OpenStudy (accessdenied):

So you have this equation now f'(x) = -cos x + e^(2x) / 2 Notice your initial conditions now include one for f'(x). f'(0) =1/2

OpenStudy (accessdenied):

f'(x) = -cos x + e^(2x)/2 + C wow i forgot the constant, bad me. That constnt is important!

OpenStudy (accessdenied):

The goal of that initial condition now is to solve for the constant. We can plug in x=0 and have f'(0) on one side, and the computed right side will become a 1-"variable" equation with C.

OpenStudy (anonymous):

so plug in 0 or 1/2??

OpenStudy (accessdenied):

f'(x) = -cos x + e^(2x) / 2 + C <-- f'(0) = 1/2 f'(0) = -cos (0) + e^(2)0)) / 2 + C ^ ^ we can compute this part this is 1/2

OpenStudy (anonymous):

so the e^0 would cancel out and what would become of -cos(0)???

OpenStudy (accessdenied):

-cos (0) You know the unit circle? The x-value of the angle 0 is the full radius. and e^(0) is equal to 1, not 0. Easy thing to overlook. :)

OpenStudy (anonymous):

oh so -1 + 1/2 +c!

OpenStudy (accessdenied):

Yes! :) We have the equality to deal with here then: 1/2 = -1 + 1/2 + C Solve for C, and we're half way there.

OpenStudy (anonymous):

1=c and now what happens??/

OpenStudy (accessdenied):

It goes back into the equation: f'(x) = -cos x + e^(2x) / 2 + 1 <--- C = 1 now. and, for the most part, we do the whole process again, starting by taking the integral of both sides.

OpenStudy (accessdenied):

But next time, we're going to use that initial condition of f(0) = 1/4, because we will then have the function f alone without any derivatives.

OpenStudy (anonymous):

so take the integral again and then plug in 0 again?

OpenStudy (accessdenied):

Pretty much, yep. If the initial condition were for some reason f(1), we'd plug in x=1 to find the constant, but this case it is made easy. to use x=0

OpenStudy (anonymous):

so sinx+e^2x +0?

OpenStudy (accessdenied):

that would be the derivative, getting back to the second derivative.

OpenStudy (anonymous):

oh right itd be -sinx +e^2x+1 /3 +x?

OpenStudy (accessdenied):

hmm.. I agree with -sin x. when we integrate e^(2x) / 2 again, that constant is kept. 1/2 times the integral of e^(2x) dx. you already took that integral before, remember what you had then?

OpenStudy (accessdenied):

\( \displaystyle \int \dfrac{e^{2x}}{2} \ dx = \frac{1}{2} \int e^{2x} \ dx \) just to illustrate the coefficient

OpenStudy (anonymous):

e^2x /2?

OpenStudy (accessdenied):

yep. that was the integral of e^(2x). and then we just multiply back the 1/2 in front: 1/2 * (e^(2x)/2) makes sense?

OpenStudy (anonymous):

can we write it as e^2x /4?

OpenStudy (accessdenied):

yes., that would be fine. so: f(x) = -sin x + e^(2x) / 4 + .... just a bit more work when we integrate a constant, it gains an x factor (a sort of reverse power rule, the derivative of x is 1.. so the integral of 1 is x) \( \int 1 dx = x + C \)

OpenStudy (anonymous):

twas what i said before! haha but without the c. so the integral is -sinx + e^2x /4 + x +c?

OpenStudy (accessdenied):

yep, that's right. sorry, saw the + 1/3 + x and wasn't quite sure. :p anyways, you have f(x) = -sin x + e^(2x) / 4 + x + C and now you can use f(0) = 1/4 to find the last constant.

OpenStudy (anonymous):

no worries! is the answer c=0???

OpenStudy (accessdenied):

Yep. So your final answer is f(x) when you plug C = 0 back in.

OpenStudy (anonymous):

its not done!!!!!!!!???? right so would i plug c=0 into the original sinx+e^2x?

OpenStudy (accessdenied):

into f(x) = -sin x + e^(2x)/4 + x + C <--- that C = 0 THEN you're done! :p

OpenStudy (accessdenied):

Basically, this f(x) is one particular solution to the original equation that satisfies those two initial conditions. If you took the derivative twice, you 'd get the original equation again.

OpenStudy (anonymous):

AH TOOK 5EVER BUT THANKS SO MUCH:-) could you help me out on more equations?

OpenStudy (accessdenied):

Uhhhh weird, my old post is there ??? Sorry, it is getting late ! (1:26 AM) I am sure there are others eager to help, though! :D

OpenStudy (accessdenied):

These problems do tend to take a while, but if you understand the process and get good practice with it, it can go a lot faster. :D

OpenStudy (anonymous):

its ok no worries! i have a test on this soon so thank u for reverything!!1